OpenStudy (anonymous):

the ac source has a peak value of 20 mv. find the total current and voltage across the si diode.

OpenStudy (anonymous):

hi do you have the circuit drawing of this one? Vrms=(Vm)/sqrt2=20mV/sqrt2=14.14mV Vdc= 2(Vm)/sqrt 2 Vdc=2(20mV)/pi= 13mV since we dont have the drawing, we assume the Resistance in series with the silicon diode is R=1k ohm, therefore using kirchoffs voltage law E= id Rd+ iR iR=14.14 mV- 13mV=1.14 mV current i= 1.14mV/ 1K ohm= 1.14 micro V if there s a circuit drawing maybe i can help you out... :D